OpenStudy (anonymous):

Solve the equation 2 over 5 (x − 2) = 4x for x. 2 over 9 9 negative 2 over 9 negative 9 over 2

OpenStudy (anonymous):

2 / 5 (x − 2) = 4x 2 (x − 2) = 20x 2x − 4 = 20x Solve the rest through simple addition and subtraction

OpenStudy (anonymous):

? it wouldnt make sense

OpenStudy (anonymous):

It does I just did it and I got one of the answers

OpenStudy (anonymous):

i ended up looking at 4=18x ? O.o

OpenStudy (anonymous):

It should be -4=18x You are still not done. You need to divide both sides by 18

OpenStudy (anonymous):

-o.22222222222222222222?

OpenStudy (anonymous):

Yes that's the answer simplified. I think you used a calculator. The answers are not simplified so if you need to know what answer it is then simplify all the answer choices. That or you can do it by hand and simplify it while keeping the fraction -4/18= x

OpenStudy (anonymous):

idk which one it is for one of the choices tho

OpenStudy (anonymous):

simplify all the answer choices then....plug the answers choices into a calculator and see which one gives you -o.22222222222222222222?